### v1

# Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory

Abstract
The evaluation of large language models (LLMs) via benchmarks is widespread, yet inconsistencies between different leaderboards and poor separability among top models raise concerns about their ability to accurately reflect authentic model capabilities. This paper provides a critical analysis of benchmark effectiveness, examining main-stream prominent LLM benchmarks using results from diverse models. We first propose a new framework for accurate and reliable estimations of item characteristics and model abilities. Specifically, we propose Pseudo-Siamese Network for Item Response Theory (PSN-IRT), an enhanced Item Response Theory framework that incorporates a rich set of item parameters within an IRT-grounded architecture. Based on PSN-IRT, we conduct extensive analysis which reveals significant and varied shortcomings in the measurement quality of current benchmarks. Furthermore, we demonstrate that leveraging PSN-IRT is able to construct smaller benchmarks while maintaining stronger alignment with human preference [^1] Code and data are openly available at https://github.com/Joe-Hall-Lee/PSN-IRT.
Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory
[…]
footnotetext: * The first two authors contributed equally.
1 Introduction
Figure 1: Illustration of weak separability and ranking inconsistencies in LLM benchmarks [^3] https://opencompass.org.cn.
As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important. Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions aimed at assessing model behavior across different tasks. However, in practice, existing benchmarks often exhibit significant limitations, prompting deeper consideration of their effectiveness McIntosh et al. (2024).
As illustrated in Figure 1, on one hand, different benchmarks often produce inconsistent leaderboards Perlitz et al. (2024), with the same model achieving substantially varying rankings across evaluations. This makes it challenging for researchers and developers to determine a model’s real performance. On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement. Given these limitations, there is a growing need for a more systematic analysis of LLM benchmarks.
In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets. Our experiment is based on Item Response Theory (IRT) Lord et al. (1968); Baker and Kim (2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^4] In the paper, an item refers to a benchmark sample.. However, traditional IRT struggles with the complexity and scale of modern datasets and its assumption of normally distributed abilities does not always hold.
To advance our objective, we propose the Pseudo-Siamese Network for IRT (PSN-IRT), a framework for comprehensive and interpretable analysis of benchmark test sets. PSN-IRT processes model identifiers and item identifiers through independent neural network pathways. These pathways respectively estimate latent model ability and a rich set of item parameters. The estimated properties are then integrated using an IRT-based formulation to predict outcome probabilities. This architectural design ensures both powerful modeling capabilities and strong theoretical interpretability.
To validate the effectiveness of PSN-IRT, we conduct extensive experiments on evaluation results from 12 LLMs across 11 benchmarks. Our results show that PSN-IRT outperforms previous approaches in both parameter estimation accuracy and reliability. Based on PSN-IRT, we perform in-depth analyses of these benchmarks using key metrics, with key findings as follows:
  1. 1.
    LLM benchmarks fail to achieve simultaneous excellence across multiple measurements.
  2. 2.
    LLM benchmarks suffer from widespread saturation and insufficient difficulty ceilings, limiting their ability to challenge and accurately evaluate advanced models.
  3. 3.
    LLM benchmarks exhibit data contamination in numerous items, compromising their reliability.
Furthermore, we find that model rankings derived from high-quality datasets selected by PSN-IRT are more consistent with human preference and offer stronger discriminability among top models.
Our contributions are summarized as follows:
  1. 1.
    We propose PSN-IRT, a benchmark analysis framework demonstrating superior estimation accuracy and reliability.
  2. 2.
    We conduct extensive experiments, using PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in may aspects.
  3. 3.
    We show that selecting items with PSN-IRT leads to model rankings that better align with human preference.
2 Background
2.1 Related Work
Researchers have explored various methods to enhance the reliability and interpretability of LLM benchmark evaluations. During test set construction, many benchmarks incorporate human annotation and leverage strong LLMs to ensure item quality, detect inconsistencies, and manage item difficulty He et al. (2024); Wei et al. (2024). Beyond these construction-phase strategies, recent studies have also introduced quantitative metrics to assess benchmark behavior post-deployment. For example, Benchmark Agreement Testing (BAT) evaluates the reliability of new benchmarks by comparing them with established ones using correlation-based measures Perlitz et al. (2024); White et al. (2025). Other work proposes novel metrics like Separability with Confidence for more nuanced benchmark analysis Li et al. (2024).
Item Response Theory (IRT) Lord et al. (1968); Baker and Kim (2004) offers a psychometric framework for finer-grained, item-level analysis. Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability Byrd and Srivastava (2022); Vania et al. (2021) and, more recently, to inspire methods for improving evaluation efficiency, for instance, through optimized item selection or adaptive testing paradigms Maia Polo et al. (2024); Zhuang et al. (2024). However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored.
Figure 2: The illustration of our proposed PSN-IRT. Separate neural networks estimate model ability ($\theta$) and item parameters ($a,b,c,d$), which are then combined via the IRT formula to predict the probability of a correct response. After that, the networks can be leveraged for estimating properties for models or items, respectively.
2.2 Preliminary: Item Response Theory
Item Response Theory (IRT) Lord et al. (1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities. Unlike Classical Test Theory (CTT) DeVellis (2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
Central to IRT is the Item Characteristic Curve (ICC), a mathematical function that describes the probability of a correct response, $P(X=1\mid\theta)$, as a function of ability $\theta$. Typically logistic, the ICC visualizes how item properties influence performance, serving as the foundation for IRT.
The simplest IRT model, the One-Parameter Logistic (1PL) model, defines the ICC as:
$P(X=1\mid\theta)=\frac{1}{1+e^{-(\theta-b)}}$ (1)
where $b$ denotes item difficulty. When $\theta=b$, the probability of the examinee generating a correct response is 0.5. The 1PL assumes all items have equal discriminability, a simplification that may not hold in complex testing scenarios.
More advanced IRT models introduce additional parameters to capture diverse item properties. However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data Tsutsumi et al. (2021).
3 Pseudo-Siamese Network for IRT
In this section, we propose the Pseudo-Siamese Network for IRT (PSN-IRT), designed to diagnose benchmark item quality by analyzing LLM responses. Specifically, one property can be inferred for each model: model-ability, and four properties can be inferred for each benchmark item: discriminability, difficulty, guessing-rate and feasibility.
Model Architecture.
The PSN-IRT architecture comprises two independent networks: a model network and an item network. The model network processes one-hot encoded LLM identifiers to estimate model-ability $\theta$, while the item network handles one-hot encoded item identifiers to produce the four IRT parameters: discriminability $a$, difficulty $b$, guessing-rate $c$, and feasibility $d$.
Each network consists of three fully connected layers with ReLU activation functions, enabling efficient and stable parameter estimation. After that, these estimated properties are then processed by a Logistic Calculation Layer. This layer operates using the Four-Parameter Logistic (4PL) model, an advanced IRT formulation that extends simpler models like the 1PL to capture nuanced behaviors:
$P(X=1\mid\theta)=c+(d-c)\cdot\frac{1}{1+e^{-a(\theta-b)}}$ (2)
where each parameter controls a different aspect of benchmark item behavior:
  • The difficulty $b$ determines the ability level where the probability changes most significantly.
  • The discriminability $a$ controls the steepness.
  • The guessing-rate $c$ captures how likely models are to succeed without full understanding.
  • The feasibility $d$ represents the maximum probability that even highly proficient models will correctly answer the item.
Training and Applications.
The PSN-IRT is trained end-to-end using the observed binary response data in the form of (Model, Item, Response, Outcome), where the binary outcome indicates the correctness of each LLM on each benchmark item. During training , PSN-IRT processes encoded pairs of LLM and item identifiers through their respective network pathways. For each pair, the PSN-IRT framework first internally estimates the specific model and item properties. These estimated properties are then passed to a Logistic Calculation Layer, with the training objective as follows:
$J(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(P_{i})+(1-y_{i})\log(1-P_{i})\right]$ (3)
where $N$ is the batch size. The training objective is to minimize the discrepancy between the model’s prediction and the observed binary outcomes. Consequently, all learnable weights within both the model and item networks are updated simultaneously, concurrently optimizing the estimated properties for both models and benchmark items.
Upon completion of training, the two networks of PSN-IRT can be used for evaluating model performance and diagnosing benchmark characteristics, respectively. Specifically, the Model Network, by processing a model’s encoding, outputs an estimate of the model ability $\theta$. Meanwhile, the Item Network, given an benchmark item’s encoding, outputs its four estimated psychometric parameters $a,b,c,d$, which can be used for an in-depth diagnosis of individual item characteristics and the overall benchmark quality.
4 Experiments
4.1 Set-up
Dataset.
To evaluate existing LLM benchmarks, we select 11 datasets that vary in domain coverage. They serve as a basis for comparing different IRT-based methods and analyzing test set properties. An overview of the datasets is provided in Table 1.
To construct training data, we conduct evaluations using OpenCompass Contributors (2023). The evaluation results are collected in the form of binary outcomes, indicating whether each model answered each item correctly. Then, we construct a unified outcome matrix by aggregating item-level results across all models and benchmarks. We split the interactions into training (60%), validation (20%), and test sets (20%). The validation set is utilized to monitor performance during model training and to implement early stopping, thereby preventing overfitting.
Benchmark Data Size Domain Metric Format
ARC-C Clark et al. (2018) 295 General EM Multiple Choice
BBH Suzgun et al. (2023) 6,511 General EM Mixed
Chinese SimpleQA He et al. (2024) 3,000 Knowledge LLM-as-a-Judge QA
GPQA Diamond Rein et al. (2024) 198 Science EM Multiple Choice
GSM8K Cobbe et al. (2021) 1,319 Math EM QA
HellaSwag Zellers et al. (2019) 10,042 General EM Multiple Choice
HumanEval Chen et al. (2021) 164 Code Pass@1 QA
MATH Hendrycks et al. (2021b) 5,000 Math EM QA
MBPP Austin et al. (2021) 500 Code Pass@1 QA
MMLU Hendrycks et al. (2021a) 14,042 General EM Multiple Choice
TheoremQA Chen et al. (2023) 800 Science EM QA
Table 1: Benchmarks used in this paper.
Model Parameter Method ACC F1 AUC Kendall Average
IRT 4PL MLE 0.7211 0.8034 0.7012 0.9697 0.7989
MCMC 0.7070 0.7811 0.7278 0.9697 0.7964
VI 0.7201 0.8015 0.6940 0.9091 0.7812
VIBO 0.7188 0.8007 0.7055 0.9697 0.7987
Deep-IRT 1PL Deep Learning 0.7969 0.8507 0.8507 0.9697 0.8670
PSN-IRT 4PL Deep Learning 0.7991 0.8520 0.8477 1.0000 0.8747
Table 2: Comparison of prediction accuracy and rank reliability across different methods.
Model.
We mainly focus on currently high-performing models, including 360GPT2-Pro [^5] https://ai.360.cn/, DeepSeek-V3 DeepSeek-AI et al. (2025), Doubao-Pro [^6] https://www.volcengine.com/product/doubao, Gemini-1.5 Team et al. (2024a), Hunyuan-Turbo [^7] https://hunyuan.tencent.com/, Moonshot-v1888https://platform.moonshot.cn/, Qwen-Plus Yang et al. (2024), and Yi-Lightning Wake et al. (2025). However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality Martínez-Plumed et al. (2019). To address this, we intentionally included several relatively weaker models, including Gemma-2B-it Team et al. (2024b), Mistral-7B-Instruct-v0.1 Jiang et al. (2023), Qwen2.5-3B-Instruct Yang et al. (2024), and Vicuna-7B-v1.3 Chiang et al. (2023).
Baseline.
We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) Hastings (1970), Variational Inference (VI) Jordan et al. (1999), and VIBO Wu et al. (2020). We also compare with Deep-IRT Tsutsumi et al. (2021), which is a neural extension of 1PL IRT that learns item-trait interactions via deep networks.
Metric.
Two quantitative metrics are included:
  • Prediction accuracy: We measure how well each method predicts whether a model answers an item correctly. These metrics include accuracy, F1 score, and ROC AUC.
  • Rank stability: We assess the consistency of model rankings by splitting the test set into two subsets. For each subset, we estimate model abilities and generate a ranking. Kendall’s $\tau$ is then computed between the two rankings.
4.2 Estimation Ability of PSN-IRT
Table 2 shows the estimation accuracy and reliability of PSN-IRT compared to traditional IRT and Deep-IRT. Both Deep-IRT and PSN-IRT significantly outperform traditional IRT in prediction accuracy. Despite its simpler Logistic output layer rooted in the IRT formula, PSN-IRT attains prediction accuracy on par with the more complex Deep-IRT, and surpasses it in rank reliability.
Given our primary objective of analyzing test sets, all methods except Deep-IRT utilized the more comprehensive 4PL structure for modeling benchmark items. However, the experimental results indicate that employing different parameter structures has a minimal impact on prediction accuracy, with detailed results available in Appendix A.3.
Dataset Difficulty Discriminability Guessing Feasibility LEH Fisher Total
Chinese SimpleQA 2 3 1 9 3 5 21
MBPP 5 6 4 8 4 4 26
MATH 4 1 2 7 7 10 27
TheoremQA 1 10 3 11 2 2 27
GPQA Diamond 3 11 6 10 1 1 30
MMLU 9 9 9 5 5 3 31
BBH 6 5 8 6 6 7 32
GSM8K 8 2 5 3 11 11 32
HellaSwag 10 7 10 2 8 6 33
HumanEval 7 4 7 4 9 9 33
ARC-C 11 8 11 1 10 8 38
Table 3: Ranks of datasets based on average item-level properties estimated by PSN-IRT. Each dataset’s score is the mean of its benchmark items’ property. Lower values indicate better performance in that property.
Figure 3: Distribution of item-level properties across 11 LLM benchmarks.
5 Benchmark Analysis with PSN-IRT
5.1 Benchmark Quality Measurements
In this section, we delve into a detailed analysis of the benchmarks based on different measurements. Specifically, we leverage the 4 item parameters (difficulty, discriminability, guessing-rate and feasibility) estimated by PSN-IRT, along with two additional metrics for evaluating item quality: Local Efficiency Headroom (LEH) Vania et al. (2021) and Fisher information Lord (1980).
LEH score assesses the potential of a test example to evaluate future progress in LLMs. It is calculated as the derivative of the item’s Item Characteristic Curve (ICC) with respect to the highest observed latent ability. A high LEH score indicates that even the top model remains far from the saturation point of the ICC.
Fisher information measures the amount of information an item provides about a model’s ability level. Items with higher Fisher information are more informative for estimating model abilities. For 4PL IRT, it is defined as:
$I(\theta)=\frac{a^{2}(P(\theta)-c)^{2}(d-P(\theta))^{2}}{(d-c)^{2}P(\theta)(1-P(\theta))},$ (4)
where $P(\theta)$ is defined as Equation 2.
Figure 4: Scatter plots showing the relationship between item difficulty and discriminability across 11 benchmarks. Each plot highlights how difficulty affects discriminability within a dataset.
5.2 Benchmark Analysis
Based on the measurements, we present an aggregate rank for each benchmark based on the sum of individual ranks in Table 3, where a lower total score indicates better overall quality. We also present the distribution of item-level properties across 11 LLM benchmarks in Figure 3. Moreover, we further illuminates the relationship between item difficulty and discriminability in Figure 4. Based on these experiment results, we present our detailed analysis as follows.
Finding 1: LLM benchmarks fail to achieve simultaneous excellence across multiple measurement properties.
As shown in Table 3, the results clearly demonstrate this lack of simultaneous excellence. While Chinese SimpleQA achieves the best overall score, no benchmark performs universally well across the individual metrics. For instance, Chinese SimpleQA itself suffers from poor feasibility with a rank of 9. This pattern of strengths being offset by significant weaknesses is common across benchmarks, which underscores the inherent challenges and trade-offs in benchmark design. Moreover, the property distributions visualized in Figure 3 further highlight this variability and the difficulty for any single benchmark to achieve balanced, high quality across all desired measurement properties.
Finding 2: Current LLM benchmarks exhibit an insufficient difficulty ceiling to challenge the most advanced models.
As shown in Table 3, difficulty is varied among different benchmarks. Datasets such as TheoremQA, Chinese SimpleQA, and GPQA Diamond lead in difficulty and continue to challenge many LLMs. In contrast, other benchmarks including ARC-C, HellaSwag, and HumanEval have become comparatively easy, where many of the items are readily solved by high-performing models, diminishing their utility for differentiating top-tier systems.
Crucially, as shown in Figure 3, even within more challenging benchmarks, items often present only moderate difficulty to elite LLMs. For instance, item difficulty scores in the IRT scaling may be as low as 1.0 for models exhibiting ability scores above 3.0. This suggests that while such benchmarks can assess average models, they are less effective at separating most advanced systems, suggesting that they are insufficient for comprehensively evaluating frontier model capabilities.
Finding 3: Generally low LEH scores across most benchmarks reveal widespread item saturation.
While Table 3 indicates that datasets like GPQA Diamond and TheoremQA achieve the highest relative ranks for average LEH, suggesting they offer more headroom than other benchmarks, the absolute LEH values in Figure 3 are often smaller than ideal. Conversely, benchmarks such as MATH and particularly GSM8K exhibit even lower average LEH scores. This signifies that current high-performing models are already approaching or have reached the performance ceiling for a substantial portion of items within these test sets. This necessitates the development of new benchmark items and datasets explicitly designed with greater headroom as models continue to evolve.
Finding 4: Outlier-high guessing-rates in many benchmark items flag risks of data contamination.
The guessing-rate reflects the chance that a model can answer a question correctly without actual understanding, typically by exploiting format-based cues or shortcuts. As shown in Figure 3, datasets like ARC-C, HellaSwag, and MMLU generally present higher guessing-rates, possibly due to their multiple-choice formats. Notably, further inspection of the item guessing-rate parameter distributions, as depicted in Figure 3, reveals that a considerable number of items across several benchmarks exhibit unusually high guessing-rates. This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data Zhuang et al. (2024).
Finding 5: Widespread low-feasibility items in scientific benchmarks reveal flawed question design.
Feasibility estimates how well a question can be answered based on the information provided. Low feasibility often arises from underspecified or overly broad questions, where even a strong model cannot determine a clear or defensible answer. As shown in Figure 3, datasets like TheoremQA and GPQA Diamond exhibit lower feasibility, likely due to complex scientific contexts or vague problem descriptions. In contrast, ARC-C and HellaSwag have high feasibility, suggesting their questions are well-formed and specific. Low feasibility more often reflects weaknesses in annotation quality or task design, and such items can distort evaluation by penalizing models for failing to resolve ambiguities Rodriguez et al. (2021).
Finding 6: Item discriminability is weakened by excessively high or low item difficulty.
As shown in Table 3, benchmarks such as MATH and GSM8K achieve high discriminability, while GPQA Diamond and TheoremQA exhibit the lowest discriminability. This limited variability in model outcomes on these latter datasets may arise from too high item difficulty, thereby reducing their capacity for precise model differentiation.
Moreover, as shown in Figure 4, item discriminability tends to be lower when item difficulty reaches extreme levels—being either too easy or too hard. Consequently, datasets that feature a balanced distribution of difficulty often exhibit higher discriminability, enabling them to more effectively distinguish between models of varying capabilities.
6 Efficient Benchmarking with PSN-IRT
The pursuit of comprehensive LLM evaluation has often led to the creation of increasingly large and diverse benchmark datasets. However, the sheer volume of benchmark items does not inherently guarantee higher quality or more insightful evaluations, and can incur significant computational costs. In this section, we explore how PSN-IRT can facilitate more efficient benchmarking by guiding the strategic curation of smaller, yet highly informative, item sets that maintain or even enhance evaluation accuracy and reliability compared to larger, undifferentiated collections.
To empirically examine this principle, we design an experiment to evaluate the effectiveness of different item selection strategies. Our methodology involve Benchmark Agreement Testing (BAT) Perlitz et al. (2024). We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena Chiang et al. (2024) and OpenCompass Arena Contributors (2023) [^9] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities.. Subsequently, various subsets of benchmark items are chosen based on different criteria. Model rankings generated using these subsets are then compared against the reference arena ranking using Kendall’s $\tau$ to measure agreement. Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models. The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess differentiation among stronger contenders.
As shown in Table 4, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking. For instance, a carefully selected subset composed of just 1000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048. This level of agreement markedly surpasses that achieved by rankings derived from using all available items or from similarly sized randomly selected subsets.
In contrast, other item selection metrics such as discriminability proves less effective for differentiating among strong models. While such items are typically adept at distinguishing between broad capability tiers (e.g., strong versus weak models), they often lack the granularity to capture nuanced performance differences at the higher end of the ability spectrum. This further verify the estimation reliability of our proposed PSN-IRT.
Metric w/ weak models w/o weak models
Variance Kendall Variance Kendall
All 0.0434 0.6444 0.0010 0.2381
Top 400
Random 0.0421 0.6444 0.0007 0.2381
Discrim. 0.1841 0.5556 0.0000 0.2381
Fisher 0.0049 0.6889 0.0033 0.7143
Top 1000
Random 0.0406 0.6444 0.0009 0.2381
Discrim. 0.1805 0.6000 0.0000 0.3813
Fisher 0.0039 0.6889 0.0030 0.9048
Table 4: Variance and Kendall’s $\tau$ for various data selection strategies with and without weak models.
7 Conclusion
In this work, we introduce PSN-IRT, a framework that combines psychometrics with neural networks to provide comprehensive and reliable analyses of LLM benchmarks. Through extensive experiments on 12 LLMs across 11 diverse datasets, we demonstrate that current benchmarks often suffer from uneven measurement properties, insufficient difficulty ceilings, item saturation, and data contamination. Moreover, we show that strategically selecting smaller, high-quality item sets using PSN-IRT enhances alignment with human preference and separability among top models.
Limitations
Our work still has some limitations: 1) The current analysis is based on a specific selection of contemporary LLMs and benchmarks. Future studies could extend our analysis to encompass a broader and more diverse range of models and datasets as they emerge. 2) PSN-IRT is designed to process binary (correct/incorrect) outcome data. This restricts its direct applicability to benchmarks that utilize graded scoring Zheng et al. (2023), for which extensions incorporating more generalized psychometric models would be necessary. 3) The item parameters estimated by PSN-IRT are currently dependent on the items being included in the training data, meaning that analyzing entirely new, unseen items typically requires retraining the model. Future iterations could explore the integration of item content features to potentially enable parameter estimation for new items without full retraining.
Appendix A Appendix
A.1 Detailed Statistics
We provide the full list of average item-level property estimates for each dataset in Table 5. These values are computed as the mean of all item parameters within a benchmark and include difficulty, discriminability, guessing-rate, feasibility, LEH score, and Fisher information. The table allows for fine-grained comparisons between benchmarks beyond their aggregated rankings.
Dataset Difficulty Discriminability Guessing Feasibility LEH Fisher
ARC-C -1.1637 1.2822 0.2621 0.9481 0.0041 0.0005
BBH -0.6033 1.3845 0.1657 0.8747 0.0057 0.0006
Chinese SimpleQA 0.2656 1.4474 0.0609 0.7088 0.0089 0.0006
GPQA Diamond 0.1902 0.8598 0.1518 0.6437 0.0142 0.0014
GSM8K -0.7899 1.5255 0.1491 0.9167 0.0038 0.0004
HellaSwag -1.0353 1.2920 0.2449 0.9304 0.0047 0.0006
HumanEval -0.6351 1.4459 0.1565 0.8957 0.0046 0.0005
MATH -0.2396 1.5793 0.0673 0.8290 0.0054 0.0004
MBPP -0.3456 1.3625 0.1345 0.8221 0.0071 0.0007
MMLU -0.8383 1.1968 0.2381 0.8921 0.0067 0.0008
TheoremQA 0.6548 1.0656 0.0888 0.5012 0.0145 0.0013
Table 5: Average item-level properties across datasets. Each value represents the mean of the corresponding metric over all benchmark items.
A.2 Case Study
Figure 5: Examples of benchmark items exhibiting high and low estimated values for key psychometric parameters, as determined by PSN-IRT. Each pair shows an item with a high parameter value alongside one with a low value for the specified characteristic.
As depicted in Figure 5, these parameters capture a wide range of item behaviors:
  • For difficulty, the figure contrasts a complex coding task with a simple factual recall question, showcasing items at opposite ends of the challenge spectrum.
  • Examples for discriminability include a quantitative reasoning problem with a high value, indicating strong differentiating power, and a common-sense inference task with a low value, suggesting limited ability to distinguish between models.
  • For guessing-rate, an alphabetical sorting task exhibits a very low estimated parameter, offering little opportunity for chance success. In contrast, a multiple-choice question concerning scientific studies shows an exceptionally high guessing parameter. This item’s content is publicly accessible online, suggesting an associated risk of data contamination.
  • For feasibility, a logical arrangement task shows high feasibility, implying clarity and high attainability for proficient models, while a domain-specific question exhibits low feasibility, suggesting potential item ambiguities.
A.3 Ablation Study
To investigate the impact of model configurations, we evaluate PSN-IRT and baseline methods across 1PL to 4PL parameters, as shown in Table 6. Among these models, the Two-Parameter Logistic (2PL) model builds upon the 1PL by incorporating an item discriminability parameter $a$:
$P(X=1\mid\theta,a,b)=\frac{1}{1+e^{-a(\theta-b)}}$ (5)
Subsequently, the Three-Parameter Logistic (3PL) model extends the 2PL by adding a guessing-rate parameter $c$:
$P(X=1\mid\theta,a,b,c)=c+(1-c)\cdot\frac{1}{1+e^{-a(\theta-b)}}$ (6)
The results demonstrate that performance exhibits no consistent pattern with increasing parameter complexity. While PSN-IRT with 4PL achieves the highest performance, simpler models occasionally outperform more complex ones in traditional IRT methods.
Additionally, we explore replacing one-hot encodings with learned embeddings and substituting the multi-layer perceptron with a graph neural network to model LLM-item interactions. However, these modifications failed to yield significant improvements, likely due to the inherent simplicity of LLM-item interaction patterns that don’t require complex relational modeling. We will explore improved input representations and model architectures as future work to enhance model robustness.
Model Method Parameter ACC F1 AUC Kendall Average
IRT MLE 1PL 0.7219 0.8035 0.6970 1.0000 0.8056
2PL 0.7172 0.7998 0.6963 0.9091 0.7806
3PL 0.7218 0.8031 0.7057 0.9697 0.8001
4PL 0.7211 0.8034 0.7012 0.9697 0.7989
MCMC 1PL 0.7168 0.8011 0.7268 0.9697 0.8036
2PL 0.7232 0.8130 0.7301 0.9697 0.8090
3PL 0.7172 0.8081 0.7297 1.0000 0.8138
4PL 0.7070 0.7811 0.7278 0.9697 0.7964
VI 1PL 0.7209 0.8024 0.6822 0.9697 0.7938
2PL 0.7198 0.8012 0.6938 1.0000 0.8037
3PL 0.7189 0.8011 0.6956 0.9697 0.7963
4PL 0.7201 0.8015 0.6940 0.9091 0.7812
VIBO 1PL 0.7007 0.7801 0.6901 0.9697 0.7852
2PL 0.7301 0.8048 0.6987 0.8788 0.7781
3PL 0.7172 0.7993 0.7048 0.9091 0.7826
4PL 0.7188 0.8007 0.7055 0.9697 0.7987
Deep-IRT Deep Learning 1PL 0.7969 0.8507 0.8507 0.9697 0.8670
PSN-IRT Deep Learning 1PL 0.7948 0.8497 0.8473 0.9091 0.8502
2PL 0.7560 0.8266 0.8259 0.9697 0.8445
3PL 0.7965 0.8510 0.8510 0.8788 0.8443
4PL 0.7991 0.8520 0.8477 1.0000 0.8747
Table 6: Comparison of prediction accuracy and rank reliability across different methods.

### v2

@@
 The evaluation of large language models (LLMs) via benchmarks is widespread, yet inconsistencies between different leaderboards and poor separability among top models raise concerns about their ability to accurately reflect authentic model capabilities.
-This paper provides a critical analysis of benchmark effectiveness, examining main-stream prominent LLM benchmarks using results from diverse models.
-We first propose a new framework for accurate and reliable estimations of item characteristics and model abilities.
-Specifically, we propose Pseudo-Siamese Network for Item Response Theory (PSN-IRT), an enhanced Item Response Theory framework that incorporates a rich set of item parameters within an IRT-grounded architecture.
-Based on PSN-IRT, we conduct extensive analysis which reveals significant and varied shortcomings in the measurement quality of current benchmarks.
-Furthermore, we demonstrate that leveraging PSN-IRT is able to construct smaller benchmarks while maintaining stronger alignment with human preference [^1] Code and data are openly available at https://github.com/Joe-Hall-Lee/PSN-IRT.
-Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory
-[…]
-footnotetext: * The first two authors contributed equally.
+This paper provides a critical analysis of benchmark effectiveness, examining mainstream prominent LLM benchmarks using results from diverse models.
+We first propose Pseudo-Siamese Network for Item Response Theory (PSN-IRT), an enhanced Item Response Theory framework that incorporates a rich set of item parameters within an IRT-grounded architecture.
+PSN-IRT can be utilized for accurate and reliable estimations of item characteristics and model abilities.
+Based on PSN-IRT, we conduct extensive analysis on 11 LLM benchmarks comprising 41,871 items, revealing significant and varied shortcomings in their measurement quality.
+Furthermore, we demonstrate that leveraging PSN-IRT is able to construct smaller benchmarks while maintaining stronger alignment with human preference.
+Codehttps://github.com/Joe-Hall-Lee/PSN-IRT
 1 Introduction
+As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important.
+Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions to assess model behavior across different tasks.
+However, in practice, existing benchmarks often exhibit significant limitations, prompting consideration of their effectiveness (McIntosh et al.
+2024).
 Figure 1: Illustration of weak separability and ranking inconsistencies in LLM benchmarks [^2] https://opencompass.org.cn.
-As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important.
-Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions aimed at assessing model behavior across different tasks.
-However, in practice, existing benchmarks often exhibit significant limitations, prompting deeper consideration of their effectiveness McIntosh et al.
-(2024).
-As illustrated in Figure 1, on one hand, different benchmarks often produce inconsistent leaderboards Perlitz et al.
-(2024), with the same model achieving substantially varying rankings across evaluations.
-This makes it challenging for researchers and developers to determine a model’s real performance.
+As illustrated in Figure 2, on one hand, different benchmarks often produce inconsistent leaderboards (Perlitz et al.
+2024).
+This issue persists even for benchmarks designed to evaluate similar underlying capabilities, leading to the same model achieving substantially varying rankings across evaluations.
 On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement.
@@
 In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets.
-Our experiment is based on Item Response Theory (IRT) Lord et al.
-(1968); Baker and Kim (2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^4] In the paper, an item refers to a benchmark sample..
-However, traditional IRT struggles with the complexity and scale of modern datasets and its assumption of normally distributed abilities does not always hold.
+Our experiment is based on Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^3] In the paper, an item refers to a benchmark sample..
+However, traditional IRT struggles with the complexity and scale of modern datasets, and its assumption of the normal distribution of abilities does not always hold.
 To advance our objective, we propose the Pseudo-Siamese Network for IRT (PSN-IRT), a framework for comprehensive and interpretable analysis of benchmark test sets.
@@
 Our results show that PSN-IRT outperforms previous approaches in both parameter estimation accuracy and reliability.
-Based on PSN-IRT, we perform in-depth analyses of these benchmarks using key metrics, with key findings as follows:
+Based on PSN-IRT, we perform in-depth analyses of these benchmarks using key metrics, with main findings as follows:
   1. 1.
@@
   1. 1.
-    We propose PSN-IRT, a benchmark analysis framework demonstrating superior estimation accuracy and reliability.
+    We propose PSN-IRT, a benchmark analysis framework with superior estimation accuracy and reliability.
   2. 2.
-    We conduct extensive experiments, using PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in may aspects.
+    We use PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in many aspects.
   3. 3.
     We show that selecting items with PSN-IRT leads to model rankings that better align with human preference.
-2 Background
-2.1 Related Work
-Researchers have explored various methods to enhance the reliability and interpretability of LLM benchmark evaluations.
-During test set construction, many benchmarks incorporate human annotation and leverage strong LLMs to ensure item quality, detect inconsistencies, and manage item difficulty He et al.
-(2024); Wei et al.
-(2024).
-Beyond these construction-phase strategies, recent studies have also introduced quantitative metrics to assess benchmark behavior post-deployment.
-For example, Benchmark Agreement Testing (BAT) evaluates the reliability of new benchmarks by comparing them with established ones using correlation-based measures Perlitz et al.
-(2024); White et al.
-(2025).
-Other work proposes novel metrics like Separability with Confidence for more nuanced benchmark analysis Li et al.
-(2024).
-Item Response Theory (IRT) Lord et al.
-(1968); Baker and Kim (2004) offers a psychometric framework for finer-grained, item-level analysis.
-Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability Byrd and Srivastava (2022); Vania et al.
-(2021) and, more recently, to inspire methods for improving evaluation efficiency, for instance, through optimized item selection or adaptive testing paradigms Maia Polo et al.
-(2024); Zhuang et al.
-(2024).
-However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored.
 Figure 2: The illustration of our proposed PSN-IRT.
@@
 After that, the networks can be leveraged for estimating properties for models or items, respectively.
+2 Background
+2.1 Related Work
+Researchers have explored various methods to evaluate and analyze the quality of LLM benchmarks.
+Recent efforts have introduced quantitative metrics that operate at a holistic, dataset level.
+For example, Benchmark Agreement Testing (BAT) assesses benchmark reliability by comparing the consistency of model rankings across different leaderboards (Perlitz et al.
+2024; White et al.
+2025).
+Other novel metrics have also been proposed for more nuanced dataset-level analysis (Li et al.
+2024).
+Delving deeper than macro-level comparisons, another significant line of work focuses on content-based analysis to ground evaluation in concrete capabilities.
+For instance, ADeLe (Zhou et al.
+2025) leverages scalable cognitive rubrics and item demand features to build highly interpretable capability profiles for AI systems.
+Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004) offers a psychometric framework for data-driven, item-level analysis.
+Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability (Vania et al.
+2021; Byrd and Srivastava 2022).
+More recently, its principles have also been used to inspire methods for improving evaluation efficiency, for instance, through optimized item selection (Maia Polo et al.
+2024) or adaptive testing paradigms (Zhuang et al.
+2024).
+However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored (Ye et al.
+2025).
 2.2 Preliminary: Item Response Theory
-Item Response Theory (IRT) Lord et al.
-(1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities.
-Unlike Classical Test Theory (CTT) DeVellis (2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
+Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities.
+Unlike Classical Test Theory (CTT) (DeVellis 2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
 Central to IRT is the Item Characteristic Curve (ICC), a mathematical function that describes the probability of a correct response, $P(X=1\mid\theta)$, as a function of ability $\theta$.
@@
 More advanced IRT models introduce additional parameters to capture diverse item properties.
-However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data Tsutsumi et al.
-(2021).
+However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data (Tsutsumi, Kinoshita, and Ueno 2021).
 3 Pseudo-Siamese Network for IRT
@@
   • The difficulty $b$ determines the ability level where the probability changes most significantly.
-  • The discriminability $a$ controls the steepness.
+  • The discriminability $a$ reflects an item’s power to differentiate between models of varying abilities.
   • The guessing-rate $c$ captures how likely models are to succeed without full understanding.
@@
 Meanwhile, the Item Network, given an benchmark item’s encoding, outputs its four estimated psychometric parameters $a,b,c,d$, which can be used for an in-depth diagnosis of individual item characteristics and the overall benchmark quality.
-4 Experiments
-4.1 Set-up
-Dataset.
+4 Efficacy of PSN-IRT
+4.1 Experimental Setup
+Datasets.
 To evaluate existing LLM benchmarks, we select 11 datasets that vary in domain coverage.
@@
 An overview of the datasets is provided in Table 1.
-To construct training data, we conduct evaluations using OpenCompass Contributors (2023).
+To construct training data, we conduct evaluations using OpenCompass (Contributors 2023).
 The evaluation results are collected in the form of binary outcomes, indicating whether each model answered each item correctly.
 Then, we construct a unified outcome matrix by aggregating item-level results across all models and benchmarks.
-We split the interactions into training (60%), validation (20%), and test sets (20%).
-The validation set is utilized to monitor performance during model training and to implement early stopping, thereby preventing overfitting.
+We split the interactions into training, validation, and test sets.
 Benchmark Data Size Domain Metric Format
-ARC-C Clark et al. (2018) 295 General EM Multiple Choice
-BBH Suzgun et al. (2023) 6,511 General EM Mixed
-Chinese SimpleQA He et al. (2024) 3,000 Knowledge LLM-as-a-Judge QA
-GPQA Diamond Rein et al. (2024) 198 Science EM Multiple Choice
-GSM8K Cobbe et al. (2021) 1,319 Math EM QA
-HellaSwag Zellers et al. (2019) 10,042 General EM Multiple Choice
-HumanEval Chen et al. (2021) 164 Code Pass@1 QA
-MATH Hendrycks et al. (2021b) 5,000 Math EM QA
-MBPP Austin et al. (2021) 500 Code Pass@1 QA
-MMLU Hendrycks et al. (2021a) 14,042 General EM Multiple Choice
-TheoremQA Chen et al. (2023) 800 Science EM QA
+ARC-C (Clark et al. 2018) 295 General EM Multiple Choice
+BBH (Suzgun et al. 2023) 6,511 General EM Mixed
+Chinese SimpleQA (He et al. 2024) 3,000 Knowledge LLM-as-a-Judge QA
+GPQA Diamond (Rein et al. 2024) 198 Science EM Multiple Choice
+GSM8K (Cobbe et al. 2021) 1,319 Math EM QA
+HellaSwag (Zellers et al. 2019) 10,042 General EM Multiple Choice
+HumanEval (Chen et al. 2021) 164 Code Pass@1 QA
+MATH (Hendrycks et al. 2021b) 5,000 Math EM QA
+MBPP (Austin et al. 2021) 500 Code Pass@1 QA
+MMLU (Hendrycks et al. 2021a) 14,042 General EM Multiple Choice
+TheoremQA (Chen et al. 2023) 800 Science EM QA
 Table 1: Benchmarks used in this paper.
-Model Parameter Method ACC F1 AUC Kendall Average
-IRT 4PL MLE 0.7211 0.8034 0.7012 0.9697 0.7989
-MCMC 0.7070 0.7811 0.7278 0.9697 0.7964
-VI 0.7201 0.8015 0.6940 0.9091 0.7812
-VIBO 0.7188 0.8007 0.7055 0.9697 0.7987
-Deep-IRT 1PL Deep Learning 0.7969 0.8507 0.8507 0.9697 0.8670
-PSN-IRT 4PL Deep Learning 0.7991 0.8520 0.8477 1.0000 0.8747
+Model Parameter Method ACC F1 AUC Kendall’s $\tau$
+IRT 4PL MLE 0.7211 0.8034 0.7012 0.9697
+MCMC 0.7070 0.7811 0.7278 0.9697
+VI 0.7201 0.8015 0.6940 0.9091
+VIBO 0.7188 0.8007 0.7055 0.9697
+Deep-IRT 1PL Deep Learning 0.7974 0.8516 0.8519 0.9697
+PSN-IRT 4PL Deep Learning 0.7998 0.8538 0.8485 1.0000
 Table 2: Comparison of prediction accuracy and rank reliability across different methods.
-Model.
-We mainly focus on currently high-performing models, including 360GPT2-Pro [^5] https://ai.360.cn/, DeepSeek-V3 DeepSeek-AI et al.
-(2025), Doubao-Pro [^6] https://www.volcengine.com/product/doubao, Gemini-1.5 Team et al.
-(2024a), Hunyuan-Turbo [^7] https://hunyuan.tencent.com/, Moonshot-v1888https://platform.moonshot.cn/, Qwen-Plus Yang et al.
-(2024), and Yi-Lightning Wake et al.
-(2025).
-However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality Martínez-Plumed et al.
-(2019).
-To address this, we intentionally included several relatively weaker models, including Gemma-2B-it Team et al.
-(2024b), Mistral-7B-Instruct-v0.1 Jiang et al.
-(2023), Qwen2.5-3B-Instruct Yang et al.
-(2024), and Vicuna-7B-v1.3 Chiang et al.
-(2023).
-Baseline.
-We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) Hastings (1970), Variational Inference (VI) Jordan et al.
-(1999), and VIBO Wu et al.
-(2020).
-We also compare with Deep-IRT Tsutsumi et al.
-(2021), which is a neural extension of 1PL IRT that learns item-trait interactions via deep networks.
-Metric.
-Two quantitative metrics are included:
-  • Prediction accuracy: We measure how well each method predicts whether a model answers an item correctly. These metrics include accuracy, F1 score, and ROC AUC.
-  • Rank stability: We assess the consistency of model rankings by splitting the test set into two subsets.
-For each subset, we estimate model abilities and generate a ranking.
-Kendall’s $\tau$ is then computed between the two rankings.
-4.2 Estimation Ability of PSN-IRT
+We mainly focus on currently high-performing models, including 360GPT2-Pro [^4] https://ai.360.cn/, DeepSeek-V3 (DeepSeek-AI et al.
+2025), Doubao-Pro [^5] https://www.volcengine.com/product/doubao, Gemini-1.5 (Team et al.
+2024a), Hunyuan-Turbo [^6] https://hunyuan.tencent.com/, Moonshot-v1777https://platform.moonshot.cn/, Qwen-Plus (Qwen et al.
+2025), and Yi-Lightning (Wake et al.
+2025).
+However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality (Mart´ınez-Plumed et al.
+2019).
+To address this, we intentionally included several relatively weaker models, including Gemma-2B-it (Team et al.
+2024b), Mistral-7B-Instruct-v0.1 (Jiang et al.
+2023), Qwen2.5-3B-Instruct (Qwen et al.
+2025), and Vicuna-7B-v1.3 (Chiang et al.
+2023).
+Baselines.
+We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) (Hastings 1970), Variational Inference (VI) (Jordan et al.
+1999), and VIBO (Wu et al.
+2020).
+We also compare with Deep-IRT (Tsutsumi, Kinoshita, and Ueno 2021), which is an extension of 1PL IRT that learns item-trait interactions via deep networks.
+Metrics.
+To assess whether the learned parameters meaningfully reflect corresponding properties, we use two quantitative metrics:
+  • Prediction accuracy: Following standard practice in educational testing (Eignor 2013), we assess how well a method predicts whether a model answers an item correctly, reporting accuracy, F1 score, and ROC AUC.
+  • Rank stability: To evaluate the reliability of model ranking induced by each method, we split the test set into two subsets, estimate model abilities separately, and compute Kendall’s $\tau$ between the resulting rankings.
+4.2 Main Results
 Table 2 shows the estimation accuracy and reliability of PSN-IRT compared to traditional IRT and Deep-IRT.
@@
 Despite its simpler Logistic output layer rooted in the IRT formula, PSN-IRT attains prediction accuracy on par with the more complex Deep-IRT, and surpasses it in rank reliability.
-Given our primary objective of analyzing test sets, all methods except Deep-IRT utilized the more comprehensive 4PL structure for modeling benchmark items.
-However, the experimental results indicate that employing different parameter structures has a minimal impact on prediction accuracy, with detailed results available in Appendix A.3.
+Given our goal to comprehensively capture diverse item characteristics, we adopt the more expressive 4PL structure for all methods except Deep-IRT.
+Experiments also confirm that PSN-IRT performs best with 4PL, with detailed results provided in the Appendix B.1.
+4.3 Ablation Study
+To assess whether more complex architectures offer benefits over our straightforward design, we conduct an ablation study.
+Specifically, we compare PSN-IRT with two variants: one replaces one-hot inputs with semantic embeddings from Instructor (Su et al.
+2023), and the other uses a GNN (Su et al.
+2022) instead of the MLP backbone [^8] Please refer to Appendix A.2 for more details..
+As shown in Table 3, both alternatives underperform our original design.
+We argue that semantic similarity between item texts does not necessarily reflect similarity in measurement characteristics.
+For the GNN variant, its neighborhood aggregation mechanism risks diluting the unique statistical signals of individual models and items, thereby hindering precise parameter estimation.
+These results suggest that in this estimation scenario, preserving the individuality of models and items is more effective than introducing architectural complexity.
+Method ACC F1 AUC Kendall
+PSN-IRT 0.7998 0.8538 0.8485 1.0000
+   + Embedding 0.7808 0.8413 0.8310 0.9394
+   + GNN 0.7928 0.8490 0.8197 0.7273
+Table 3: Ablation study on PSN-IRT’s input representation and network architecture.
+Furthermore, since PSN-IRT estimates parameters that are specific to the training data, requiring retraining to analyze new items, we also investigate the stability of these estimates when the item pool changes.
+To simulate this, We conduct an experiment by independently training separate PSN-IRT models on random subsets of the items.
+We then compute the Pearson correlation between the item parameters estimated from these subsets and those estimated from the full dataset.
+As shown in Table 4, the results reveal a high degree of correlation across all parameters.
+This indicates that PSN-IRT learns highly consistent intrinsic properties for the items regardless of the specific data sample, demonstrating the stability of parameter estimation.
+Data Difficulty Discrim. Guessing Feasibility
+30% 0.9009 0.8186 0.8274 0.9025
+50% 0.7437 0.7835 0.8776 0.9330
+70% 0.9519 0.7414 0.8320 0.9442
+Table 4: Pearson correlation coefficients of item parameters estimated on random subsets of the data versus those estimated on the full dataset.
+All correlations are statistically significant ($p<0.0001$).
 Dataset Difficulty Discriminability Guessing Feasibility LEH Fisher Total
@@
 In this section, we delve into a detailed analysis of the benchmarks based on different measurements.
-Specifically, we leverage the 4 item parameters (difficulty, discriminability, guessing-rate and feasibility) estimated by PSN-IRT, along with two additional metrics for evaluating item quality: Local Efficiency Headroom (LEH) Vania et al.
-(2021) and Fisher information Lord (1980).
+Specifically, we leverage the 4 item parameters (difficulty, discriminability, guessing-rate and feasibility) estimated by PSN-IRT, along with two additional metrics for evaluating item quality: Local Efficiency Headroom (LEH) (Vania et al.
+2021) and Fisher information (Lord 1980).
 LEH score assesses the potential of a test example to evaluate future progress in LLMs.
@@
 where $P(\theta)$ is defined as Equation 2.
-Figure 4: Scatter plots showing the relationship between item difficulty and discriminability across 11 benchmarks. Each plot highlights how difficulty affects discriminability within a dataset.
 5.2 Benchmark Analysis
-Based on the measurements, we present an aggregate rank for each benchmark based on the sum of individual ranks in Table 3, where a lower total score indicates better overall quality.
-We also present the distribution of item-level properties across 11 LLM benchmarks in Figure 3.
-Moreover, we further illuminates the relationship between item difficulty and discriminability in Figure 4.
-Based on these experiment results, we present our detailed analysis as follows.
+Based on the measurements, we provide an aggregate rank for each benchmark, calculated from the sum of individual ranks in Table 5, where a lower total score indicates better overall quality.
+Additionally, Figure 3 visualizes the distribution of item-level properties across the 11 LLM benchmarks.
+Building on these experimental results, our detailed analysis is presented as follows.
 Finding 1: LLM benchmarks fail to achieve simultaneous excellence across multiple measurement properties.
@@
 Notably, further inspection of the item guessing-rate parameter distributions, as depicted in Figure 3, reveals that a considerable number of items across several benchmarks exhibit unusually high guessing-rates.
-This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data Zhuang et al.
-(2024).
+This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data (Zhuang et al.
+2024).
 Finding 5: Widespread low-feasibility items in scientific benchmarks reveal flawed question design.
 Feasibility estimates how well a question can be answered based on the information provided.
-Low feasibility often arises from underspecified or overly broad questions, where even a strong model cannot determine a clear or defensible answer.
+Low feasibility often arises from underspecified or overly broad questions, where even a strong model cannot determine a clear answer.
 As shown in Figure 3, datasets like TheoremQA and GPQA Diamond exhibit lower feasibility, likely due to complex scientific contexts or vague problem descriptions.
 In contrast, ARC-C and HellaSwag have high feasibility, suggesting their questions are well-formed and specific.
-Low feasibility more often reflects weaknesses in annotation quality or task design, and such items can distort evaluation by penalizing models for failing to resolve ambiguities Rodriguez et al.
-(2021).
-Finding 6: Item discriminability is weakened by excessively high or low item difficulty.
-As shown in Table 3, benchmarks such as MATH and GSM8K achieve high discriminability, while GPQA Diamond and TheoremQA exhibit the lowest discriminability.
-This limited variability in model outcomes on these latter datasets may arise from too high item difficulty, thereby reducing their capacity for precise model differentiation.
-Moreover, as shown in Figure 4, item discriminability tends to be lower when item difficulty reaches extreme levels—being either too easy or too hard.
-Consequently, datasets that feature a balanced distribution of difficulty often exhibit higher discriminability, enabling them to more effectively distinguish between models of varying capabilities.
+Low feasibility more often reflects weaknesses in annotation quality or task design, and such items can distort evaluation by penalizing models for failing to resolve ambiguities (Rodriguez et al.
+2021).
 6 Efficient Benchmarking with PSN-IRT
-The pursuit of comprehensive LLM evaluation has often led to the creation of increasingly large and diverse benchmark datasets.
-However, the sheer volume of benchmark items does not inherently guarantee higher quality or more insightful evaluations, and can incur significant computational costs.
-In this section, we explore how PSN-IRT can facilitate more efficient benchmarking by guiding the strategic curation of smaller, yet highly informative, item sets that maintain or even enhance evaluation accuracy and reliability compared to larger, undifferentiated collections.
-To empirically examine this principle, we design an experiment to evaluate the effectiveness of different item selection strategies.
-Our methodology involve Benchmark Agreement Testing (BAT) Perlitz et al.
-(2024).
-We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena Chiang et al.
-(2024) and OpenCompass Arena Contributors (2023) [^9] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities..
+While the pursuit of comprehensive LLM evaluation has led to increasingly large benchmarks, sheer item volume neither guarantees quality nor comes without significant computational cost.
+This section, therefore, explores how PSN-IRT can facilitate more efficient benchmarking.
+We investigate whether strategically curating smaller, highly informative item sets can yield model comparisons that maintain or even surpass the accuracy and reliability of larger, undifferentiated collections.
+Experimental Setup.
+To evaluate the effectiveness of different item selection strategies, we design an experiment to evaluate the effectiveness of different item selection strategies.
+Our methodology involve Benchmark Agreement Testing (BAT) (Perlitz et al.
+2024).
+We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena (Chiang et al.
+2024) and OpenCompass Arena (Contributors 2023) [^9] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities..
 Subsequently, various subsets of benchmark items are chosen based on different criteria.
@@
 Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models.
-The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess differentiation among stronger contenders.
-As shown in Table 4, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking.
-For instance, a carefully selected subset composed of just 1000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048.
+The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess separability among stronger contenders.
+Results.
+As shown in Table 6, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking.For instance, a carefully selected subset composed of just 1,000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048.
 This level of agreement markedly surpasses that achieved by rankings derived from using all available items or from similarly sized randomly selected subsets.
-In contrast, other item selection metrics such as discriminability proves less effective for differentiating among strong models.
-While such items are typically adept at distinguishing between broad capability tiers (e.g., strong versus weak models), they often lack the granularity to capture nuanced performance differences at the higher end of the ability spectrum.
-This further verify the estimation reliability of our proposed PSN-IRT.
-Metric w/ weak models w/o weak models
+Method w/ weak models w/o weak models
 Variance Kendall Variance Kendall
@@
 Random 0.0421 0.6444 0.0007 0.2381
+Clustering 0.0308 0.6000 0.0058 0.1429
 Discrim. 0.1841 0.5556 0.0000 0.2381
@@
 Random 0.0406 0.6444 0.0009 0.2381
+Clustering 0.0166 0.6444 0.0041 0.2381
 Discrim. 0.1805 0.6000 0.0000 0.3813
@@
 Table 6: Variance and Kendall’s $\tau$ for various data selection strategies with and without weak models.
+In contrast, other item selection approaches are less effective for differentiating among strong models.
+For instance, selecting items based on discriminability, results in zero score variance when weak models are excluded, indicating it merely separates broad capability tiers rather than providing granularity among top-performers.
+Separately, a strategy of clustering items based on their success/failure vectors (Pacchiardi, Cheke, and Hernández-Orallo 2024) also fails to improve ranking correlation, even when it increases score separability.
 7 Conclusion
@@
 Moreover, we show that strategically selecting smaller, high-quality item sets using PSN-IRT enhances alignment with human preference and separability among top models.
-Limitations
-Our work still has some limitations: 1) The current analysis is based on a specific selection of contemporary LLMs and benchmarks.
-Future studies could extend our analysis to encompass a broader and more diverse range of models and datasets as they emerge.
-2) PSN-IRT is designed to process binary (correct/incorrect) outcome data.
-This restricts its direct applicability to benchmarks that utilize graded scoring Zheng et al.
-(2023), for which extensions incorporating more generalized psychometric models would be necessary.
-3) The item parameters estimated by PSN-IRT are currently dependent on the items being included in the training data, meaning that analyzing entirely new, unseen items typically requires retraining the model.
-Future iterations could explore the integration of item content features to potentially enable parameter estimation for new items without full retraining.
-Appendix A Appendix
-A.1 Detailed Statistics
+Appendix A Implementation Details
+A.1 Estimation Ability of PSN-IRT
+To evaluate the estimation ability of PSN-IRT against baselines, we first construct a unified response matrix by aggregating item-level results across all models and benchmarks.
+We then split these model-item interactions into training (60%), validation (20%), and test sets (20%).
+The model is trained on the training set using the Adam optimizer with a learning rate of 0.003 and a weight decay of 0.0001, and a batch size of 512.
+To ensure robust model selection and prevent overfitting, we employ an early stopping strategy with a patience of 5 epochs, monitored based on the F1 score on the validation set.
+For the subsequent benchmark analysis presented in Section 5, where the goal is to obtain the most stable parameter estimates for all items, we train the PSN-IRT model on the entire unified dataset.
+In this phase, the model is trained for a fixed 30 epochs using the same hyperparameters.
+All training is conducted on a single NVIDIA A800-80G GPU.
+A.2 Ablation Study
+This section provides additional implementation details for the ablation study variants mentioned in Section 4.3.
+Semantic Embedding Variant.
+To explore the utility of item content, we substitute one-hot encoded item identifiers with dense semantic embeddings.
+These embeddings are generated for the text of each benchmark item using the pre-trained Instructor-XL [^10] https://huggingface.co/hkunlp/instructor-xl.
+They then serve as the direct input to the item network, replacing the one-hot encoding pathway while the rest of the PSN-IRT architecture remained unchanged.
+However, these variants fail to yield significant improvements. We will explore improved input representations and model architectures as future work to enhance model robustness.
+Graph Neural Network (GNN) Variant.
+To investigate the role of network architecture, we implemented a custom GNN-based encoder.
+This encoder updates a model’s initial embedding by aggregating information from all interactions in the training graph.
+For each interaction between a model $s$ and an item $q$ with a binary outcome $rel\in\{0,1\}$, a cognitive degree $\mathbf{l}_{s,q}$ is first computed:
+$\mathbf{l}_{s,q}=\text{Linear}_{\text{gate}}([\mathbf{e}_{q},\mathbf{W}_{rel}\cdot rel])$ (5)
+where $\mathbf{e}_{q}$ is the item’s initial embedding, $\mathbf{W}_{rel}$ is a trainable weight matrix that projects the scalar response $rel$ into the embedding space, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation.
+This combined vector is then processed by a single linear layer ($\text{Linear}_{\text{gate}}$) to form the cognitive degree.
+Subsequently, a cognitive state $\mathbf{c}_{s,q}$ is formed through an element-wise product:
+$\mathbf{c}_{s,q}=\mathbf{e}_{q}\odot\mathbf{l}_{s,q}$ (6)
+The updated embedding for model $s$, denoted $\mathbf{e}^{\prime}_{s}$, is obtained by performing an attention-weighted aggregation of the cognitive states from all items $\mathcal{N}_{s}$ it has interacted with:
+$\mathbf{e}^{\prime}_{s}=\mathbf{e}_{s}+\sum_{q_{i}\in\mathcal{N}_{s}}\alpha_{s,q_{i}}\cdot\mathbf{c}_{s,q_{i}}$ (7)
+where the attention weights $\alpha_{s,q_{i}}$ are calculated via a softmax over scores.
+These scores are derived by applying a linear layer and a LeakyReLU activation to a concatenation of the model’s embedding $\mathbf{e}_{s}$ and each cognitive state $\mathbf{c}_{s,q_{i}}$.
+A symmetric process is used to update item embeddings.
+However, both variants fail to yield performance gains. We leave the exploration of more specialized architectures and feature representations as future work.
+Appendix B Additional Analysis
+B.1 Performance Comparison across IRT Parmaters
+To investigate the impact of model configurations, we evaluate PSN-IRT and baseline methods across 1PL to 4PL parameters, as shown in Table 7.
+Among these models, the Two-Parameter Logistic (2PL) model builds upon the 1PL by incorporating an item discriminability parameter $a$:
+$P(X=1\mid\theta,a,b)=\frac{1}{1+e^{-a(\theta-b)}}$ (8)
+Subsequently, the Three-Parameter Logistic (3PL) model extends the 2PL by adding a guessing-rate parameter $c$:
+$P(X=1\mid\theta,a,b,c)=c+(1-c)\cdot\frac{1}{1+e^{-a(\theta-b)}}$ (9)
+The results demonstrate that performance exhibits no consistent pattern with increasing parameter complexity.
+While PSN-IRT with 4PL achieves the highest performance, simpler models occasionally outperform more complex ones in traditional IRT methods.
+Model Method Parameter ACC F1 AUC Kendall’s $\tau$
+IRT MLE 1PL 0.7219 0.8035 0.6970 1.0000
+2PL 0.7172 0.7998 0.6963 0.9091
+3PL 0.7218 0.8031 0.7057 0.9697
+4PL 0.7211 0.8034 0.7012 0.9697
+MCMC 1PL 0.7168 0.8011 0.7268 0.9697
+2PL 0.7232 0.8130 0.7301 0.9697
+3PL 0.7172 0.8081 0.7297 1.0000
+4PL 0.7070 0.7811 0.7278 0.9697
+VI 1PL 0.7209 0.8024 0.6822 0.9697
+2PL 0.7198 0.8012 0.6938 1.0000
+3PL 0.7189 0.8011 0.6956 0.9697
+4PL 0.7201 0.8015 0.6940 0.9091
+VIBO 1PL 0.7007 0.7801 0.6901 0.9697
+2PL 0.7301 0.8048 0.6987 0.8788
+3PL 0.7172 0.7993 0.7048 0.9091
+4PL 0.7188 0.8007 0.7055 0.9697
+Deep-IRT Deep Learning 1PL 0.7974 0.8516 0.8519 0.9697
+PSN-IRT Deep Learning 1PL 0.7948 0.8497 0.8473 0.9091
+2PL 0.7560 0.8266 0.8259 0.9697
+3PL 0.7965 0.8510 0.8510 0.8788
+4PL 0.7998 0.8538 0.8485 1.0000
+Table 7: Comparison of prediction accuracy and rank reliability across different methods.
+B.2 Detailed Benchmark Statistics
 We provide the full list of average item-level property estimates for each dataset in Table 8.
@@
 Table 8: Average item-level properties across datasets. Each value represents the mean of the corresponding metric over all benchmark items.
-A.2 Case Study
+B.3 Relationship Between Item Difficulty and Discriminability
+Our analysis reveals a notable relationship between item difficulty and discriminability.
+As illustrated by the scatter plots in Figure 4, item discriminability tends to be lower when item difficulty reaches extreme levels—that is, when items are either very easy or very hard.
+Consequently, datasets that feature a more balanced distribution of difficulty often exhibit higher overall discriminability, enabling them to more effectively distinguish between models of varying capabilities.
+This pattern is also reflected in the dataset-level averages presented in Table 5 in the main text.
+For instance, benchmarks such as MATH and GSM8K, which contain a range of difficulties, achieve high average discriminability.
+Conversely, datasets like GPQA Diamond and TheoremQA, which are characterized by a concentration of very difficult items, exhibit the lowest average discriminability.
+This suggests that their capacity for precise model differentiation is reduced, in part, due to this difficulty-discriminability trade-off.
+These observations underscore the importance of balancing difficulty when designing test items to ensure a benchmark is not only challenging but also highly discriminative.
+Figure 4: Scatter plots showing the relationship between item difficulty and discriminability across 11 benchmarks. Each plot highlights how difficulty affects discriminability within a dataset.
+B.4 Case Study
 Figure 5: Examples of benchmark items exhibiting high and low estimated values for key psychometric parameters, as determined by PSN-IRT.
@@
   • For difficulty, the figure contrasts a complex coding task with a simple factual recall question, showcasing items at opposite ends of the challenge spectrum.
-  • Examples for discriminability include a quantitative reasoning problem with a high value, indicating strong differentiating power, and a common-sense inference task with a low value, suggesting limited ability to distinguish between models.
+  • For discriminability, the examples include a quantitative reasoning problem with a high value, indicating strong differentiating power, and a common-sense inference task with a low value, suggesting limited ability to distinguish between models.
   • For guessing-rate, an alphabetical sorting task exhibits a very low estimated parameter, offering little opportunity for chance success.
@@
   • For feasibility, a logical arrangement task shows high feasibility, implying clarity and high attainability for proficient models, while a domain-specific question exhibits low feasibility, suggesting potential item ambiguities.
-A.3 Ablation Study
-To investigate the impact of model configurations, we evaluate PSN-IRT and baseline methods across 1PL to 4PL parameters, as shown in Table 6.
-Among these models, the Two-Parameter Logistic (2PL) model builds upon the 1PL by incorporating an item discriminability parameter $a$:
-$P(X=1\mid\theta,a,b)=\frac{1}{1+e^{-a(\theta-b)}}$ (5)
-Subsequently, the Three-Parameter Logistic (3PL) model extends the 2PL by adding a guessing-rate parameter $c$:
-$P(X=1\mid\theta,a,b,c)=c+(1-c)\cdot\frac{1}{1+e^{-a(\theta-b)}}$ (6)
-The results demonstrate that performance exhibits no consistent pattern with increasing parameter complexity.
-While PSN-IRT with 4PL achieves the highest performance, simpler models occasionally outperform more complex ones in traditional IRT methods.
-Additionally, we explore replacing one-hot encodings with learned embeddings and substituting the multi-layer perceptron with a graph neural network to model LLM-item interactions.
-However, these modifications failed to yield significant improvements, likely due to the inherent simplicity of LLM-item interaction patterns that don’t require complex relational modeling.
-We will explore improved input representations and model architectures as future work to enhance model robustness.
-Model Method Parameter ACC F1 AUC Kendall Average
-IRT MLE 1PL 0.7219 0.8035 0.6970 1.0000 0.8056
-2PL 0.7172 0.7998 0.6963 0.9091 0.7806
-3PL 0.7218 0.8031 0.7057 0.9697 0.8001
-4PL 0.7211 0.8034 0.7012 0.9697 0.7989
-MCMC 1PL 0.7168 0.8011 0.7268 0.9697 0.8036
-2PL 0.7232 0.8130 0.7301 0.9697 0.8090
-3PL 0.7172 0.8081 0.7297 1.0000 0.8138
-4PL 0.7070 0.7811 0.7278 0.9697 0.7964
-VI 1PL 0.7209 0.8024 0.6822 0.9697 0.7938
-2PL 0.7198 0.8012 0.6938 1.0000 0.8037
-3PL 0.7189 0.8011 0.6956 0.9697 0.7963
-4PL 0.7201 0.8015 0.6940 0.9091 0.7812
-VIBO 1PL 0.7007 0.7801 0.6901 0.9697 0.7852
-2PL 0.7301 0.8048 0.6987 0.8788 0.7781
-3PL 0.7172 0.7993 0.7048 0.9091 0.7826
-4PL 0.7188 0.8007 0.7055 0.9697 0.7987
-Deep-IRT Deep Learning 1PL 0.7969 0.8507 0.8507 0.9697 0.8670
-PSN-IRT Deep Learning 1PL 0.7948 0.8497 0.8473 0.9091 0.8502
-2PL 0.7560 0.8266 0.8259 0.9697 0.8445
-3PL 0.7965 0.8510 0.8510 0.8788 0.8443
-4PL 0.7991 0.8520 0.8477 1.0000 0.8747
-Table 6: Comparison of prediction accuracy and rank reliability across different methods.

### v3

@@
 1 Introduction
-As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important.
-Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions to assess model behavior across different tasks.
+As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important (Chang et al.
+2024; Zhou et al.
+2024; Huang et al.
+2025).
+Currently, the performance of LLMs is primarily evaluated through various benchmarks (Wang et al.
+2024), which are comprehensive test suites consisting of carefully designed questions to assess model behavior across different tasks.
 However, in practice, existing benchmarks often exhibit significant limitations, prompting consideration of their effectiveness (McIntosh et al.
@@
 Figure 1: Illustration of weak separability and ranking inconsistencies in LLM benchmarks [^2] https://opencompass.org.cn.
-As illustrated in Figure 2, on one hand, different benchmarks often produce inconsistent leaderboards (Perlitz et al.
+As illustrated in Figure 2, on one hand, even benchmarks that are designed to measure similar underlying capabilities often produce inconsistent leaderboards, leading to substantial ranking variations for the same model (Perlitz et al.
 2024).
-This issue persists even for benchmarks designed to evaluate similar underlying capabilities, leading to the same model achieving substantially varying rankings across evaluations.
 On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement.
@@
 In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets.
-Our experiment is based on Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^3] In the paper, an item refers to a benchmark sample..
+Our experiment is based on Item Response Theory (IRT) (Lord et al.
+1968; Baker and Kim 2004), a psychometric framework widely used in educational assessment to analyze item properties such as difficulty by modeling item response against examinee ability [^3] In the paper, an item refers to a benchmark sample..
 However, traditional IRT struggles with the complexity and scale of modern datasets, and its assumption of the normal distribution of abilities does not always hold.
@@
   2. 2.
-    We use PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in many aspects.
+    We use PSN-IRT for an in-depth analysis of mainstream benchmarks and find that current LLM benchmarks present deficiencies in many aspects.
   3. 3.
@@
 Figure 2: The illustration of our proposed PSN-IRT.
-Separate neural networks estimate model ability ($\theta$) and item parameters ($a,b,c,d$), which are then combined via the IRT formula to predict the probability of a correct response.
+Separate neural networks estimate model-ability ($\theta$) and item parameters ($a,b,c,d$), which are then combined via the IRT formula to predict the probability of a correct response.
 After that, the networks can be leveraged for estimating properties for models or items, respectively.
@@
 2025) leverages scalable cognitive rubrics and item demand features to build highly interpretable capability profiles for AI systems.
-Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004) offers a psychometric framework for data-driven, item-level analysis.
-Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability (Vania et al.
+Item Response Theory (IRT) (Lord et al.
+1968; Baker and Kim 2004) offers a psychometric framework for data-driven, item-level analysis.
+Within natural language processing (NLP), IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability (Vania et al.
 2021; Byrd and Srivastava 2022).
@@
 2024) or adaptive testing paradigms (Zhuang et al.
-2024).
+2025; Truong et al.
+2025).
 However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored (Ye et al.
@@
 2.2 Preliminary: Item Response Theory
-Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities.
+Item Response Theory (IRT) (Lord et al.
+1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities.
 Unlike Classical Test Theory (CTT) (DeVellis 2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
@@
 More advanced IRT models introduce additional parameters to capture diverse item properties.
-However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data (Tsutsumi, Kinoshita, and Ueno 2021).
+However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data (Tsutsumi et al.
+2021).
 3 Pseudo-Siamese Network for IRT
@@
   • The feasibility $d$ represents the maximum probability that even highly proficient models will correctly answer the item.
+Notably, PSN-IRT follows standard IRT assumptions such as unidimensionality and monotonicity, which generally hold for LLM benchmarking (Kipnis et al. 2025; Truong et al. 2025).
 Training and Applications.
@@
 Upon completion of training, the two networks of PSN-IRT can be used for evaluating model performance and diagnosing benchmark characteristics, respectively.
-Specifically, the Model Network, by processing a model’s encoding, outputs an estimate of the model ability $\theta$.
-Meanwhile, the Item Network, given an benchmark item’s encoding, outputs its four estimated psychometric parameters $a,b,c,d$, which can be used for an in-depth diagnosis of individual item characteristics and the overall benchmark quality.
+Specifically, the model network, by processing a model’s encoding, outputs an estimate of the model-ability $\theta$.
+Meanwhile, the item network, given a benchmark item’s encoding, outputs its four estimated psychometric parameters $a,b,c,d$, which can be used for an in-depth diagnosis of individual item characteristics and the overall benchmark quality.
 4 Efficacy of PSN-IRT
@@
 Table 2: Comparison of prediction accuracy and rank reliability across different methods.
-We mainly focus on currently high-performing models, including 360GPT2-Pro [^4] https://ai.360.cn/, DeepSeek-V3 (DeepSeek-AI et al.
-2025), Doubao-Pro [^5] https://www.volcengine.com/product/doubao, Gemini-1.5 (Team et al.
-2024a), Hunyuan-Turbo [^6] https://hunyuan.tencent.com/, Moonshot-v1777https://platform.moonshot.cn/, Qwen-Plus (Qwen et al.
-2025), and Yi-Lightning (Wake et al.
-2025).
-However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality (Mart´ınez-Plumed et al.
+We mainly focus on currently high-performing models, including 360GPT2-Pro [^4] https://ai.360.cn/, DeepSeek-V3555https://platform.deepseek.com/, Doubao-pro [^6] https://www.volcengine.com/product/doubao, Gemini-1.5777https://ai.google.dev/, Hunyuan-Turbo [^8] https://hunyuan.tencent.com/, Moonshot-v1999https://platform.moonshot.cn/, Qwen-Plus [^10] https://qwen.ai/apiplatform, and Yi-Lightning [^11] https://platform.lingyiwanwu.com/.
+However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality (Martínez-Plumed et al.
 2019).
 To address this, we intentionally included several relatively weaker models, including Gemma-2B-it (Team et al.
-2024b), Mistral-7B-Instruct-v0.1 (Jiang et al.
+2024), Mistral-7B-Instruct-v0.1 (Jiang et al.
 2023), Qwen2.5-3B-Instruct (Qwen et al.
@@
 2020).
-We also compare with Deep-IRT (Tsutsumi, Kinoshita, and Ueno 2021), which is an extension of 1PL IRT that learns item-trait interactions via deep networks.
+We also compare with Deep-IRT (Tsutsumi et al.
+2021), which is an extension of 1PL IRT that learns item-trait interactions via deep networks.
 Metrics.
@@
 Given our goal to comprehensively capture diverse item characteristics, we adopt the more expressive 4PL structure for all methods except Deep-IRT.
-Experiments also confirm that PSN-IRT performs best with 4PL, with detailed results provided in the Appendix B.1.
+Experiments also confirm that PSN-IRT performs best with 4PL.
 4.3 Ablation Study
@@
 2023), and the other uses a GNN (Su et al.
-2022) instead of the MLP backbone [^8] Please refer to Appendix A.2 for more details..
+2022) instead of the MLP backbone.
 As shown in Table 3, both alternatives underperform our original design.
@@
 Dataset Difficulty Discriminability Guessing Feasibility LEH Fisher Total
-Chinese SimpleQA 2 3 1 9 3 5 21
-MBPP 5 6 4 8 4 4 26
-MATH 4 1 2 7 7 10 27
-TheoremQA 1 10 3 11 2 2 27
-GPQA Diamond 3 11 6 10 1 1 30
-MMLU 9 9 9 5 5 3 31
-BBH 6 5 8 6 6 7 32
-GSM8K 8 2 5 3 11 11 32
-HellaSwag 10 7 10 2 8 6 33
-HumanEval 7 4 7 4 9 9 33
-ARC-C 11 8 11 1 10 8 38
+Chinese SimpleQA 2 3 1 9 3 5 23
+TheoremQA 1 10 3 11 2 2 29
+MBPP 5 6 4 8 4 4 31
+MATH 4 1 2 7 7 10 31
+GPQA Diamond 3 11 6 10 1 1 32
+BBH 6 5 8 6 6 7 38
+MMLU 9 9 9 5 5 3 40
+GSM8K 8 2 5 3 11 11 40
+HumanEval 7 4 7 4 9 9 40
+HellaSwag 10 7 10 2 8 6 43
+ARC-C 11 8 11 1 10 8 49
 Table 5: Ranks of datasets based on average item-level properties estimated by PSN-IRT.
-Each dataset’s score is the mean of its benchmark items’ property.
-Lower values indicate better performance in that property.
+Each rank is obtained by averaging the property over all items in a dataset and then sorting these averages; smaller rank values correspond to better performance.
 Figure 3: Distribution of item-level properties across 11 LLM benchmarks.
@@
 5.2 Benchmark Analysis
-Based on the measurements, we provide an aggregate rank for each benchmark, calculated from the sum of individual ranks in Table 5, where a lower total score indicates better overall quality.
+Based on the measurements, we provide an aggregate rank for each benchmark, calculated from the sum of individual ranks in Table 5, where a lower total rank value indicates better overall quality.
 Additionally, Figure 3 visualizes the distribution of item-level properties across the 11 LLM benchmarks.
@@
 Datasets such as TheoremQA, Chinese SimpleQA, and GPQA Diamond lead in difficulty and continue to challenge many LLMs.
-In contrast, other benchmarks including ARC-C, HellaSwag, and HumanEval have become comparatively easy, where many of the items are readily solved by high-performing models, diminishing their utility for differentiating top-tier systems.
-Crucially, as shown in Figure 3, even within more challenging benchmarks, items often present only moderate difficulty to elite LLMs.
-For instance, item difficulty scores in the IRT scaling may be as low as 1.0 for models exhibiting ability scores above 3.0.
-This suggests that while such benchmarks can assess average models, they are less effective at separating most advanced systems, suggesting that they are insufficient for comprehensively evaluating frontier model capabilities.
+In contrast, other benchmarks including ARC-C, HellaSwag, and MMLU have become comparatively easy, where many of the items are readily solved by high-performing models, diminishing their utility for differentiating top-tier systems.
+Crucially, as shown in Figure 3, even the most challenging items in existing benchmarks exhibit relatively low difficulty levels.
+For instance, on the IRT scale, the highest item difficulty values rarely exceed 1.0, while top models such as DeepSeek-V3 have estimated abilities well above 3.0.
+This indicates that even the hardest questions are significantly below the capability level of elite LLMs, highlighting a lack of sufficient challenge for frontier capabilities.
 Finding 3: Generally low LEH scores across most benchmarks reveal widespread item saturation.
 While Table 5 indicates that datasets like GPQA Diamond and TheoremQA achieve the highest relative ranks for average LEH, suggesting they offer more headroom than other benchmarks, the absolute LEH values in Figure 3 are often smaller than ideal.
-Conversely, benchmarks such as MATH and particularly GSM8K exhibit even lower average LEH scores.
+Conversely, benchmarks such as GSM8K and ARC-C exhibit lower average LEH scores.
 This signifies that current high-performing models are already approaching or have reached the performance ceiling for a substantial portion of items within these test sets.
@@
 This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data (Zhuang et al.
-2024).
+2025).
 Finding 5: Widespread low-feasibility items in scientific benchmarks reveal flawed question design.
@@
 To evaluate the effectiveness of different item selection strategies, we design an experiment to evaluate the effectiveness of different item selection strategies.
-Our methodology involve Benchmark Agreement Testing (BAT) (Perlitz et al.
+Our methodology involves Benchmark Agreement Testing (BAT) (Perlitz et al.
 2024).
@@
 Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models.
-The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess separability among stronger contenders.
+The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess separability among stronger contenders, with weak models introduced in Section 4.1.
 Results.
-As shown in Table 6, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking.For instance, a carefully selected subset composed of just 1,000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048.
+As shown in Table 6, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking.
+For instance, a carefully selected subset composed of just 1,000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048.
 This level of agreement markedly surpasses that achieved by rankings derived from using all available items or from similarly sized randomly selected subsets.
@@
 Table 6: Variance and Kendall’s $\tau$ for various data selection strategies with and without weak models.
-In contrast, other item selection approaches are less effective for differentiating among strong models.
-For instance, selecting items based on discriminability, results in zero score variance when weak models are excluded, indicating it merely separates broad capability tiers rather than providing granularity among top-performers.
-Separately, a strategy of clustering items based on their success/failure vectors (Pacchiardi, Cheke, and Hernández-Orallo 2024) also fails to improve ranking correlation, even when it increases score separability.
+In contrast, other item selection approaches are less effective for distinguishing strong models.
+For instance, selecting items based on discriminability yields zero score variance when weak models are excluded, indicating it merely separates broad capability tiers rather than providing granularity among top-performers.
+Separately, clustering items based on their success/failure vectors (Pacchiardi et al.
+2024) also fails to improve ranking correlation, even when it increases score separability.
 7 Conclusion
-In this work, we introduce PSN-IRT, a framework that combines psychometrics with neural networks to provide comprehensive and reliable analyses of LLM benchmarks.
+In this work, we introduce PSN‑IRT, a framework that combines psychometrics with neural networks to provide comprehensive and reliable analyses of LLM benchmarks.
 Through extensive experiments on 12 LLMs across 11 diverse datasets, we demonstrate that current benchmarks often suffer from uneven measurement properties, insufficient difficulty ceilings, item saturation, and data contamination.
-Moreover, we show that strategically selecting smaller, high-quality item sets using PSN-IRT enhances alignment with human preference and separability among top models.
+Moreover, we show that strategically selecting smaller, high‑quality item sets using PSN‑IRT enhances alignment with human preference and separability among top models.
+Acknowledgments
+This work is supported by National Natural Science Foundation of China (62276077, 62376075, 62376076), Department of Science and Technology of Heilongjiang (Grant No.
+ZY04JD04), and the Key Laboratory of Computing Power Network and Information Security, Ministry of Education (Grant No.
+2023ZD027).
 Appendix A Implementation Details
@@
 $\mathbf{l}_{s,q}=\text{Linear}_{\text{gate}}([\mathbf{e}_{q},\mathbf{W}_{rel}\cdot rel])$ (5)
-where $\mathbf{e}_{q}$ is the item’s initial embedding, $\mathbf{W}_{rel}$ is a trainable weight matrix that projects the scalar response $rel$ into the embedding space, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation.
+where $\mathbf{e}_{q}$ is the item’s initial embedding, $\mathbf{W}_{rel}$ is a trainable weight matrix that projects the scalar response relrel into the embedding space, and $\cdot$$\cdot$] denotes concatenation.
 This combined vector is then processed by a single linear layer ($\text{Linear}_{\text{gate}}$) to form the cognitive degree.
@@
 These observations underscore the importance of balancing difficulty when designing test items to ensure a benchmark is not only challenging but also highly discriminative.
-Figure 4: Scatter plots showing the relationship between item difficulty and discriminability across 11 benchmarks. Each plot highlights how difficulty affects discriminability within a dataset.
 B.4 Case Study
-Figure 5: Examples of benchmark items exhibiting high and low estimated values for key psychometric parameters, as determined by PSN-IRT.
-Each pair shows an item with a high parameter value alongside one with a low value for the specified characteristic.
 As depicted in Figure 5, these parameters capture a wide range of item behaviors:
@@
   • For feasibility, a logical arrangement task shows high feasibility, implying clarity and high attainability for proficient models, while a domain-specific question exhibits low feasibility, suggesting potential item ambiguities.
+Figure 4: Scatter plots showing the relationship between item difficulty and discriminability across 11 benchmarks. Each plot highlights how difficulty affects discriminability within a dataset.
+Figure 5: Examples of benchmark items exhibiting high and low estimated values for key psychometric parameters, as determined by PSN-IRT.
+Each pair shows an item with a high parameter value alongside one with a low value for the specified characteristic.